# Tuning the band topology and topological Hall effect in skyrmion crystals via the spin-orbit coupling

### v1

@@ III Topological Hall Effect with Rashba SOC
Figure 8: THC in the presence of the Rashba SOC term for the lowest subband. (a) The green line shows the quantized Chern number for the completely filled lowest subband ($C_{1}$) as a function of $\alpha_{R}$, which switches between zero and ±$\pm 1$, leading to a quantized THC, $\sigma_{xy}=-e^{2}/\hbar\times C_{1}$. The red line shows the THC for the half-filled lowest subband indicating that it can be continuously tuned including a sign reversal at the cross over value $\alpha_{R}^{c}$. (b)-(d) Berry curvatures $\Omega^{z}(\mathbb{k})$ in the Brillouin zone for three different values of $\alpha_{R}$ (corresponding to the three black dots in (a)) showing positive (red) and negative (blue) regions of $\Omega^{z}(\mathbb{k})$. The blue/red arrows show the anomalous velocities $e\hbar^{-1}\ \mathbb{E}\times\Omega^{z}(\mathbb{k})$ for the blue/red regions (color coded). The THC, which is the sum of $\Omega^{z}(\mathbb{k})$ over the occupied states, becomes zero for $\alpha_{R}^{c}\approx-0.27$ (Figure (c)) as the positive Berry curvature part cancels the negative part. The circles indicate the Fermi surface in each case.
@@ IV Topological Hall Effect with partially filled subbands
The computed $\sigma$ with the lowest subband half filled is shown in Fig. 8(a) as a function of $\alpha_{R}$ for the Néel skyrmion, which shows a change of sign as $\alpha_{R}$ varies through the crossover point $\alpha_{R}\approx-0.27$. This leads to the remarkable result that an applied electric field can reverse the polarity of the Hall current by tuning $\alpha_{R}$.
@@
Table 1: THC in the presence of Dresselhaus and Weyl SOC, expressed in terms of the THC for the Rashba SOC, with $\alpha$ being the SOC strength in each case. As the Table shows, the THC for all cases can be expressed in terms of just four conductivities, viz., $\sigma_{1,0}\ (\alpha)$, $\sigma_{-1,0}\ (\alpha)$, $\sigma_{1,\pi/2}\ (\alpha)$, and $\sigma_{-1,\pi/2}\ (\alpha)$, where $\sigma_{m,\gamma}\ (\alpha)\equiv\sigma_{m,\gamma}^{R}\ (\alpha)$ is the THC with Rashba SOC. For example, the table shows that $\sigma^{W}_{1,0}(\alpha)=\sigma^{R}_{1,\pi/2}(-\alpha)$, which is the first result listed under Weyl. The four independent conductivities have been plotted in Fig. 9 for the case when the lowest subband is half-filled. However, the correspondence is valid for any filling, whether we have completely filled or partially filled subbands.
@@
Vorticity Helicity $\sigma_{m,\gamma}^{R}(\alpha)$ $\sigma_{m,\gamma}^{D}(\alpha)$ $\sigma_{m,\gamma}^{W}(\alpha)$
skyrmion $m=1$ Néel $\gamma=0$ $\sigma_{1,0}(\alpha)$ $-\sigma_{-1,\pi/2}(\alpha)$ $\sigma_{1,\pi/2}(-\alpha)$
$\gamma=\pi$ $\sigma_{1,0}(-\alpha)$ $-\sigma_{-1,\pi/2}(-\alpha)$ $\sigma_{1,\pi/2}(\alpha)$
Bloch $\gamma=\pi/2$ $\sigma_{1,\pi/2}(\alpha)$ $-\sigma_{-1,0}(\alpha)$ $\sigma_{1,0}(\alpha)$
$\gamma=-\pi/2$ $\sigma_{1,\pi/2}(-\alpha)$ $-\sigma_{-1,0}(-\alpha)$ $\sigma_{1,0}(-\alpha)$
antiskyrmion $m=-1$ Néel $\gamma=0$ $\sigma_{-1,0}(\alpha)$ $-\sigma_{1,\pi/2}(\alpha)$ $\sigma_{-1,\pi/2}(-\alpha)$
$\gamma=\pi$ $\sigma_{-1,0}(-\alpha)$ $-\sigma_{1,\pi/2}(-\alpha)$ $\sigma_{-1,\pi/2}(\alpha)$
Bloch $\gamma=\pi/2$ $\sigma_{-1,\pi/2}(\alpha)$ $-\sigma_{1,0}(\alpha)$ $\sigma_{-1,0}(\alpha)$
$\gamma=-\pi/2$ $\sigma_{-1,\pi/2}(-\alpha)$ $-\sigma_{1,0}(-\alpha)$ $\sigma_{-1,0}(-\alpha)$
@@ V Dresselhaus and Weyl SOC
Apart from the Rashba SOC, one could also have a Dresselhaus or a Weyl type of SOC. While the Rashba SOC arises at the solid surface due to the surface inversion asymmetry and can be induced by an applied electric field, the Dresselhaus SOC arises due to the bulk inversion asymmetry. However, it has been demonstrated that the Dresselhaus SOC can be introduced and tuned in the 2D systems via interface engineering [57]. Furthermore, it has been recently pointed out that depending on the crystalline symmetry of the underlying material, it is possible to have Rashba, Dresselhaus, or Weyl type SOC, even at the $\Gamma$ point of the Brillouin zone[60]. In view of this, for the sake of completeness, we also consider the effect of the Dresselhaus and Weyl type SOC. The SOC terms for these cases are written in momentum space as $\mathcal{H}_{D/W}^{\rm SOC}=\alpha_{D/W}(\sigma_{x}k_{x}\mp\sigma_{y}k_{y})$ with the corresponding lattice version in real space $\mathcal{H}_{D/W}^{\rm SOC}=i\alpha_{D/W}\sum_{\langle ij\rangle}c_{i}^{\dagger}(\sigma_{x}r^{x}_{ij}\mp\sigma_{y}r^{y}_{ij})c_{j}+H.c.,$ where the (+)-(+)- ( + ) sign corresponds to the Dresselhaus (Weyl) SOC. The full Hamiltonians corresponding to Eq. (3) for the Rashba case read
@@
As for the Rashba case, the net magnetic field seen by the electron may be determined by taking an infinitesimal square loop around a spatial point and computing the flux through it from the Peierls phase accumulated by the electron while traveling around the loop. The results are:
$B^{D}_{SK}(r,\alpha)$ $=$ $\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}+\frac{\alpha_{R}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}-\frac{1}{r}\sin\frac{\pi r}{\lambda})\sin(2\alpha+\gamma),\ \ \ \ {\rm(Dresselhaus/skyrmion)}$ (14)
$B^{D}_{AS}(r)$ $=$ $-\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}+\frac{\alpha_{R}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}+\frac{1}{r}\sin\frac{\pi r}{\lambda})\sin\gamma,\ \ \ \ {\rm(Dresselhaus/antiskyrmion)}$ (15)
$B^{W}_{SK}(r)$ $=$ $\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}-\frac{\alpha_{R}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}+\frac{1}{r}\sin\frac{\pi r}{\lambda})\sin\gamma,\hskip 71.13188pt{\rm(Weyl/skyrmion)}$ (16)
$B^{W}_{AS}(r,\alpha)$ $=$ $-\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}-\frac{\alpha_{R}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}-\frac{1}{r}\sin\frac{\pi r}{\lambda})\sin(\gamma-{2\alpha}),\ \ \ \ {\rm(Weyl/antiskyrmion)}$ (17)
@@
It turns out that the Hamiltonian for the three cases of SOC (Rashba SOC, Dresselhaus SOC, or Weyl SOC) can be mapped into one another for the skyrmion types considered in the paper. As a result, the various physical quantities, in particular, the emergent magnetic fields as well as the THC are related to one another. The THC’s can be expressed, as summarized in Table 1, in terms of just four quantities, viz., $\sigma_{1,0}\ (\alpha)$, $\sigma_{-1,0}$, $\sigma_{1,\pi/2}$, and $\sigma_{-1,\pi/2}$, where $\sigma_{m,\gamma}$ is the THC in the presence of the Rashba SOC, with $m,\gamma$ being the vorticity and the helicity, respectively. These relations can be understood by noting the following points:
(1) For the THC in two cases to be the same, one must check that the emergent magnetic fields are the same, which is a necessary condition. However, this is not sufficient, since the effective hopping magnitude between two lattice sites can still be different even though the magnetic fields are the same. As an example, considering the THC $\sigma_{1,\pi/2}$ for the Rashba SOC case (green curve in Fig. 9), the magnetic field is independent of the Rashba SOC strength $\alpha_{R}$ due to the $\cos\gamma=0$ factor (see Eq. (5)). Nevertheless, the THC does change with $\alpha_{R}$ because the hopping amplitudes between two sites are functions of both $t$ and $\alpha_{R}$.
(2) The Hamiltonians possess certain symmetries. It is easy to see that in all three cases $\mathcal{H}_{R}$, $\mathcal{H}_{D}$, and $\mathcal{H}_{W}$ (Eqs. (3), (12), and (13)), when the sign of the SOC strength $\alpha$ is flipped and the helicity $m$ is changed by $\pi$ simultaneously, the Hamiltonian remains unchanged. This leads to identical band structures and all other electronic properties including transport, so that we have the equality: $\sigma_{m,\gamma}^{X}(\alpha)=\sigma_{m,\gamma\pm\pi}^{X}(-\alpha)$, where $X=R,D$, or $W$.
@@
(3) The Weyl Hamiltonian for a certain skyrmion type turns out to be the same as the Rashba Hamiltonian for certain other skyrmion type. For example, if we compare the Rashba Hamiltonian for the Néel skyrmion ($m=1,\gamma=0$) and the Weyl Hamiltonian for the Bloch skyrmion ($m=1,\gamma=\pi/2$), it is easily shown from Eqs. (3) and (13) that the two Hamiltonians are identical (see Supplementary Materials for details). Thus, we have $\sigma_{1,\pi/2}^{W}(\alpha)=\sigma_{1,0}^{R}(\alpha)$. There are several other such equalities, which have been used to construct Table 1. These have been explicitly verified by computing the THC’s for both cases.
(4) The Rashba and Dresselhaus Hamiltonians with certain skyrmion types are related to each other via a complex conjugate hopping. For instance, considering the two cases: (i) Rashba Hamiltonian for the Néel skyrmion ($m=1,\gamma=0$) and (ii) Dresselhaus Hamiltonian for the Bloch antiskyrmion ($m=-1,\gamma=\pi/2$), it is easily shown from Eqs. (3) and (12) that all terms of the two Hamiltonians are the same except that the coefficients are complex conjugate of each other as illustrated in Figure 10 ($h_{ij}d_{j}^{\dagger}d_{i}$ in one case vs. $h_{ij}^{*}d_{j}^{\dagger}d_{i}$ for the other case, mathematical expressions are given in Supplementary Materials). This means that the effective magnetic fields that the electron sees as it hops from one site to another are simply of the opposite signs. This in turn implies that the two THCs are negative of each other, because if two magnetic fields are applied to the same system in opposite directions, the sum of the two Hall conductivities must add up to zero. This leads to the equality $\sigma_{-1,\pi/2}^{D}(\alpha)=-\sigma_{1,0}^{R}(\alpha)$. Similarly, several other correspondences between the Rashba and Dresselhaus conductivities are found, which are used to produce Table 1.

### v2

@@ III Topological Hall Effect with Rashba SOC
Figure 9: THC in the presence of the Rashba SOC term for the lowest subband. (a) The green line shows the quantized Chern number for the completely filled lowest subband ($C_{1}$) as a function of $\alpha_{R}$, which switches between zero and $\pm 1$, leading to a quantized THC, $\sigma_{xy}=-e^{2}/\hbar\times C_{1}$. The red line shows the THC for the half-filled lowest subband indicating that it can be continuously tuned including a sign reversal at the crossover value $\alpha_{R}^{c}$. (b)-(d) Berry curvatures $\Omega^{z}(\mathbb{k})$ in the Brillouin zone for three different values of $\alpha_{R}$ (corresponding to the three black dots in (a)) showing positive (red) and negative (blue) regions of $\Omega^{z}(\mathbb{k})$. The blue/red arrows show the anomalous velocities $e\hbar^{-1}\ \mathbb{E}\times\Omega^{z}(\mathbb{k})$ for the blue/red regions (color coded). The THC, which is the sum of $\Omega^{z}(\mathbb{k})$ over the occupied states, becomes zero for $\alpha_{R}^{c}\approx-0.27$ (Figure (c)) as the positive Berry curvature part cancels the negative part. The circles indicate the Fermi surface in each case.
@@ IV Topological Hall Effect with partially filled subbands
The computed $\sigma$ with the lowest subband half filled is shown in Figure 9(a) as a function of $\alpha_{R}$ for the Néel skyrmion, which shows a change of sign as $\alpha_{R}$ varies through the crossover point $\alpha_{R}\approx-0.27$. This leads to the remarkable result that an applied electric field can reverse the polarity of the Hall current by tuning $\alpha_{R}$.
@@
Table 1: THC in the presence of Dresselhaus and Weyl SOC, expressed in terms of the THC for the Rashba SOC, with $\bar{\alpha}$ being the SOC strength in each case. As the Table shows, the THC for all cases can be expressed in terms of just four conductivities, viz., $\sigma_{1,0}\ (\bar{\alpha})$, $\sigma_{-1,0}\ (\bar{\alpha})$, $\sigma_{1,\pi/2}\ (\bar{\alpha})$, and $\sigma_{-1,\pi/2}\ (\bar{\alpha})$, where $\sigma_{m,\gamma}\ (\bar{\alpha})\equiv\sigma_{m,\gamma}^{R}\ (\bar{\alpha})$ is the THC with Rashba SOC. For example, the table shows that $\sigma^{W}_{1,0}(\bar{\alpha})=\sigma^{R}_{1,\pi/2}(-\bar{\alpha})$, which is the first result listed under Weyl. The four independent conductivities have been plotted in Figure 10 for the case when the lowest subband is half-filled. However, the correspondence is valid for any filling, whether we have completely filled or partially filled subbands.
@@
Vorticity Helicity $\sigma_{m,\gamma}^{R}(\bar{\alpha})$ $\sigma_{m,\gamma}^{D}(\bar{\alpha})$ $\sigma_{m,\gamma}^{W}(\bar{\alpha})$
skyrmion $m=1$ Néel $\gamma=0$ $\sigma_{1,0}(\bar{\alpha})$ $\sigma_{-1,\pi/2}(\bar{\alpha})$ $\sigma_{1,\pi/2}(-\bar{\alpha})$
$\gamma=\pi$ $\sigma_{1,0}(-\bar{\alpha})$ $\sigma_{-1,\pi/2}(-\bar{\alpha})$ $\sigma_{1,\pi/2}(\bar{\alpha})$
Bloch $\gamma=\pi/2$ $\sigma_{1,\pi/2}(\bar{\alpha})$ $\sigma_{-1,0}(\bar{\alpha})$ $\sigma_{1,0}(\bar{\alpha})$
$\gamma=-\pi/2$ $\sigma_{1,\pi/2}(-\bar{\alpha})$ $\sigma_{-1,0}(-\bar{\alpha})$ $\sigma_{1,0}(-\bar{\alpha})$
antiskyrmion m=1m=-1 Néel $\gamma=0$ $\sigma_{-1,0}(\bar{\alpha})$ $\sigma_{1,\pi/2}(\bar{\alpha})$ $\sigma_{-1,\pi/2}(-\bar{\alpha})$
$\gamma=\pi$ $\sigma_{-1,0}(-\bar{\alpha})$ $\sigma_{1,\pi/2}(-\bar{\alpha})$ $\sigma_{-1,\pi/2}(\bar{\alpha})$
Bloch $\gamma=\pi/2$ $\sigma_{-1,\pi/2}(\bar{\alpha})$ $\sigma_{1,0}(\bar{\alpha})$ $\sigma_{-1,0}(\bar{\alpha})$
$\gamma=-\pi/2$ $\sigma_{-1,\pi/2}(-\bar{\alpha})$ $\sigma_{1,0}(-\bar{\alpha})$ $\sigma_{-1,0}(-\bar{\alpha})$
@@ V Dresselhaus and Weyl SOC
Apart from the Rashba SOC, one could also have a Dresselhaus or a Weyl type of SOC. While the Rashba SOC arises at the solid surface due to the surface inversion asymmetry and can be induced by an applied electric field, the Dresselhaus SOC arises due to the bulk inversion asymmetry. However, it has been demonstrated that the Dresselhaus SOC can be introduced and tuned in the 2D systems via interface engineering [55]. Furthermore, it has been recently pointed out that depending on the crystalline symmetry of the underlying material, it is possible to have Rashba, Dresselhaus, or Weyl type SOC, even at the $\Gamma$ point of the Brillouin zone[58]. In view of this, for the sake of completeness, we also consider the effect of the Dresselhaus and Weyl type SOC. The SOC terms for these cases are written in momentum space as $\mathcal{H}_{D/W}^{\rm SOC}=\alpha_{D/W}(\sigma_{x}k_{x}\mp\sigma_{y}k_{y})$ with the corresponding lattice version in real space
$\mathcal{H}_{D/W}^{\rm SOC}=i\alpha_{D/W}\sum_{\langle ij\rangle}c_{i}^{\dagger}(\sigma_{x}r^{x}_{ij}\mp\sigma_{y}r^{y}_{ij})c_{j}+H.c.,$ (12)
where the (+)-(+) sign corresponds to the Dresselhaus (Weyl) SOC. The corresponding full Hamiltonians, similar to Eq. (3) for the Rashba case, are given by
@@
As for the Rashba case, the net magnetic field seen by the electron may be determined by taking an infinitesimal square loop around a spatial point and computing the flux through it from the Peierls phase accumulated by the electron while travelling around the loop. The results are:
$B^{D}_{SK}(r,\alpha)$ $=$ $\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}+\frac{\alpha_{D}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}-\frac{1}{r}\sin\frac{\pi r}{\lambda})\sin(2\alpha+\gamma),\ \ \ \ {\rm(Dresselhaus/skyrmion)}$ (15)
$B^{D}_{AS}(r)$ $=$ $-\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}+\frac{\alpha_{D}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}+\frac{1}{r}\sin\frac{\pi r}{\lambda})\sin\gamma,\ \ \ \ \ \ {\rm(Dresselhaus/antiskyrmion)}$ (16)
$B^{W}_{SK}(r)$ $=$ $\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}-\frac{\alpha_{W}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}+\frac{1}{r}\sin\frac{\pi r}{\lambda})\sin\gamma,\hskip 71.13188pt{\rm(Weyl/skyrmion)}$ (17)
$B^{W}_{AS}(r,\alpha)$ $=$ $-\frac{\pi}{2\lambda r}\sin\frac{\pi r}{\lambda}-\frac{\alpha_{W}}{t}(\frac{\pi}{\lambda}\cos{\frac{\pi r}{\lambda}}-\frac{1}{r}\sin\frac{\pi r}{\lambda})\sin(\gamma-{2\alpha}),\ \ \ \ \ {\rm(Weyl/antiskyrmion)}$ (18)
@@
We note that in all cases, the magnetic field expressions can be written as sum of two magnetic fields. The first one, arising out of the skyrmion spin texture, can be expressed as the well-known triple spin product[6], while the second one due to the Rashba/Dresselhaus/Weyl SOC can be expressed as spin derivatives. The corresponding magnetic fields, in units of $\Phi_{0}/(2\pi)$, are
$B~=-\frac{1}{2}\vec{S}\cdot(\partial_{x}\vec{S}\times\partial_{y}\vec{S}),~~~~~~~~{\rm(Sk/AS)}$ (19)
$B_{R}=-\frac{\alpha_{R}}{t}(\partial_{x}S_{x}+\partial_{y}S_{y}),~~~~~~~~{\rm(Rashba)}$ (20)
$B_{D}=~\frac{\alpha_{D}}{t}(\partial_{x}S_{y}+\partial_{y}S_{x}),~~~{\rm(Dresselhaus)}$ (21)
$B_{W}=-\frac{\alpha_{W}}{t}(\partial_{x}S_{y}-\partial_{y}S_{x}).~~~~~~~~{\rm(Weyl)}$ (22)
The magnetic field expressions for the Rashba and Dresselhaus cases, Eqs. (20) and (21), were obtained earlier[59], which agree with our expressions except for an overall sign for the Dresselhaus case. However, the correctness of our sign is confirmed from the fact that the relations between the THC’s presented in Table I, which we have derived analytically from the mapping of the Hamiltonians and independently checked from direct numerical calculations of the THC, would not agree otherwise. Eqs. (19 - 22), in the polar coordinate form, are provided in section III of the Supplemental Material [48].
Mapping of the Hamiltonians. It turns out that the Hamiltonian for the three cases of SOC (Rashba SOC, Dresselhaus SOC, or Weyl SOC) can be mapped into one another for the skyrmion types considered in the paper. As a result, the various physical quantities, in particular, the emergent magnetic fields as well as the THC are related to one another. The THC’s can be expressed, as summarized in Table 1, in terms of just four quantities, viz., $\sigma_{1,0}\ (\bar{\alpha})$, $\sigma_{-1,0}\ (\bar{\alpha})$, $\sigma_{1,\pi/2}\ (\bar{\alpha})$, and $\sigma_{-1,\pi/2}\ (\bar{\alpha})$, where $\sigma_{m,\gamma}\ (\bar{\alpha})$ is the THC in the presence of the Rashba SOC of strength $\bar{\alpha}$, and $m$ and $\gamma$ are the vorticity and the helicity, respectively. These relations can be understood by noting the following points:
(1) For the THC in two cases to be the same, one must check that the emergent magnetic fields are the same, which is a necessary condition. However, this is not sufficient, since the effective hopping magnitude between two lattice sites can still be different even though the magnetic fields are the same. As an example, considering the THC $\sigma_{1,\pi/2}$ for the Rashba SOC case (green curve in Figure 10), the magnetic field is independent of the Rashba SOC strength $\alpha_{R}$ due to the $\cos\gamma=0$ factor (see Eq. (5)). Nevertheless, the THC does change with $\alpha_{R}$ because the hopping amplitudes between two sites are functions of both $t$ and $\alpha_{R}$.
(2) The Hamiltonians possess certain symmetries. It is easy to see that in all three cases $\mathcal{H}_{R}$, $\mathcal{H}_{D}$, and $\mathcal{H}_{W}$ (Eqs. (3), (13), and (14)), when the sign of the SOC strength $\alpha$ is flipped and the helicity $\gamma$ is changed by $\pi$ simultaneously, the Hamiltonian remains unchanged. This leads to identical band structures and all other electronic properties including transport, so that we have the equality: $\sigma_{m,\gamma}^{X}(\bar{\alpha})=\sigma_{m,\gamma\pm\pi}^{X}(-\bar{\alpha})$, where $X=R,D$, or $W$.
@@
(3) The Weyl Hamiltonian for a certain skyrmion type turns out to be the same as the Rashba Hamiltonian for certain other skyrmion type. For example, if we compare the Rashba Hamiltonian for the Néel skyrmion ($m=1,\gamma=0$) and the Weyl Hamiltonian for the Bloch skyrmion ($m=1,\gamma=\pi/2$), it is easily shown from Eqs. (3) and (14) that the two Hamiltonians are identical [48]. Thus, we have $\sigma_{1,\pi/2}^{W}(\bar{\alpha})=\sigma_{1,0}^{R}(\bar{\alpha})$. There are several other such equalities, which have been used to construct Table 1. These relations have been explicitly verified by computing the THC’s for the various cases as well.
(4) Regarding the Rashba and Dresselhaus Hamiltonians, they turn out to be related via a complex conjugate hopping. For instance, considering the two cases: (i) Rashba Hamiltonian for the Néel skyrmion ($m=1,\gamma=0$) and (ii) Dresselhaus Hamiltonian for the Bloch antiskyrmion ($m=-1,\gamma=\pi/2$), it is easily shown from Eqs. (3) and (13) that all terms of the two Hamiltonians are the same except that the coefficients are complex conjugate of each other as illustrated in Figure 11 ($h_{ij}d_{j}^{\dagger}d_{i}$ in one case vs. $h_{ij}^{*}d_{j}^{\dagger}d_{i}$ for the other case; mathematical expressions are given in Supplementary Materials, Section IV [48]). This means that the effective magnetic fields that the electron sees as it hops from one site to another are simply of the opposite signs. This in turn implies that the two THCs are negative of each other, because if two magnetic fields are applied to the same system in opposite directions, the sum of the two Hall conductivities must add up to zero. This leads to the equality $\sigma_{-1,\pi/2}^{D}(\bar{\alpha})=-\sigma_{1,0}^{R}(\bar{\alpha})$. Similarly, several other correspondences between the Rashba and Dresselhaus conductivities are found, which are used to produce Table 1.